Geometry.
The length of a rectangle is 2 in. more than twice its width. If the
perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by checkley71(8403)

L=2+2W
2L+2W=34
2(2+2W)+2W=34
4+4W+2W=34
6W=34-4
6W=30
W=30/6
W=5
THEN L=2+2*5
L=2+10
L=12
PROOF
2*5+2*12=34
10+24=34
34=34